# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks

### v2

@@ 4.1 Qualitative results on synthetic data
Data Metric Function-space priors Weight-space priors GP-based
@@
4.2 Quantitative results on real-world data
We study the fidelity of GFSVI compared to the true GP-posterior on real data, and evaluate GFSVI on regression, classification, and out-of-distribution detection tasks. In all tables, we bold the highest score and any score whose error bar (standard error) overlaps with the highest score’s error bar.
Ocean current modeling
Figure 3: Results for the ocean current modeling experiment. We report the mean velocity vectors and the squared errors of compared methods. Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
We consider the problem of modeling ocean currents in the Gulf of Mexico to measure the extend to which GFSVI can incorporate knowledge specified via a GP prior on real-world data. Specifically, we follow the setup by Shalashilin, (2024) and Cinquin et al., (2024), and use the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) to estimate ocean currents from $20$ two-dimensional velocity vectors collected from drifter buoys. We embed physical properties of fluid motions into the GP prior and to the neural networks by applying the Helmholtz decomposition (Berlinghieri et al.,, 2023; Cinquin et al.,, 2024). We compare our GFSVI to a GP and to TFSVI. More details can be found in section A.3.2.
Table 3: Results for the ocean current modeling task. We find that incorporating prior knowledge via a Gaussian process prior in GFSVI improves score over weight-space priors in TFSVI.
Dataset GFSVI (ours) TFSVI GP
Log-like. -6.627 $\pm$ 0.753 -22.651 $\pm$ 2.947 -0.507 $\pm$ 0.000
MSE -0.021 $\pm$ 0.002 -0.034 $\pm$ 0.003 -0.013 $\pm$ 0.000
Results are shown in table 3 and in figure 3. We find that incorporating prior knowledge via an informative GP prior in GFSVI improves expected log-likelihood (Log-like.) and mean squared-error (MSE) over weight-space priors in TFSVI. However, for this relatively small dataset, exact GP inference is still possible, and it outperforms both BNN-based predictions. This suggests that the physically motivated GP describes the true fluid dynamics well enough that the additional inductive bias introduced by a neural network hurts performance rather than helping it. In the following, we consider experiments with larger datasets (making exact GP inference computationally infeasible in many cases), and where structural prior knowledge in function space exists but is not derived from strict laws of nature.
Regression
We assess the predictive performance of GFSVI on data sets from the UCI repository (Dua and Graff,, 2017) (described in table 6 in the Appendix). Table 1, and Table 7 in the Appendix, show expected log-likelihood and mean squared error, respectively. We perform 5-fold cross validation and report means and standard errors across the test folds. We also rank the methods for each dataset and report the mean rank of each method across all datasets. See section A.3.3 for more details. We find that GFSVI performs competitively compared to baselines and obtains the best mean rank for both metrics, matching the top performing methods on nearly all datasets. In particular, we find that using GP priors in the linearized BNN with GFSVI yields improvements over the weight-space priors used in TFSVI and that GFSVI performs slightly better than FVI.
@@ 4.2 Quantitative results on real-world data
Dataset Boston Concrete Energy Wine Yacht Mean rank
Classification
We further evaluate the classification performance of our method on the MNIST (LeCun et al.,, 2010) and FashionMNIST (Xiao et al.,, 2017) image data sets. We fit the models on a random subset of 90% of the training set, use the remaining 10% as validation data, and evaluate on the provided test split. We repeat with 5 different random seeds and report the mean and standard error of the expected log-likelihood, accuracy and expected calibration error (ECE) of the predictive distribution’s mean in table 2. For GFSVI, FVI, and TFSVI, we tested measurement points from both a random (rnd) distribution $\rho({\bm{x}})$ and from kmnist (see Algorithm 1 and section A.3.4). We find that GFSVI performs competitively on MNIST, exceeding the expected log-likelihood and accuracy of top-scoring baselines (see table 2) and similarly to best baselines on FashionMNIST. Additionally, GFSVI yields well-calibrated models with low ECE.
Out-of-distribution detection
We further evaluate our method by testing if its epistemic uncertainty is predictive of out-of-distribution (OOD) data. We consider two settings: (i) with tabular data and a Gaussian likelihood following Malinin et al., (2020), and (ii) with image data and a categorical likelihood following Osawa et al., (2019). We report the accuracy of classifying OOD from in-distribution (ID) data using a (learned) threshold on the predictive uncertainty. Additional details are provided in section A.3.5. In setting (i) (table 4), we find that GFSVI performs competitively and obtains the highest mean rank. Likewise, in setting (ii) (table 8 in the Appendix), GFSVI strongly outperforms all baselines when using the kmnist measurement point distribution $\rho({\bm{x}})$. We find that with high-dimensional image data, the choice of measurement point distribution highly influences OOD detection accuracy. We provide more details on this point in Appendix A.4.4. In both settings, we find that using GP priors with GFSVI rather than weight-space priors with TFSVI is beneficial, and that GFSVI improves over FVI.
Variational measure evaluation
table 5 evaluates our inference method by comparing samples drawn from the exact posterior (where computationally feasible) with the approximate posterior obtained with our method (GFSVI). More details are provided in section A.3.6. We find that GFSVI approximates the exact posterior more accurately that FVI, obtaining a higher mean rank, but worse than the gold standard sparse GP, which demonstrates to be most accurate.
@@ 5 Related work
In this section, we review related work on function-space variational inference with neural networks, and on approximating functions-space measures with weight-space priors.
Function-space inference with neural networks
Prior work on function-space VI in BNNs has addressed issues (i) and (ii) mentioned in Section 2.1. Sun et al., (2019) address (i) (intractable variational posterior in function space) by using implicit score function estimators, and (ii) (intractable KL divergence) by replacing the supremum with an expectation. Rudner et al., 2022a ; Rudner et al., 2022b address (i) by using a linearized BNN (Khan et al.,, 2020; Immer et al.,, 2021; Maddox et al.,, 2021), and (ii) by replacing the supremum with a maximum over a finite set of samples. Other work abandons approximating the posterior over neural network weights altogether and instead uses a BNN only to specify a prior (Ma et al.,, 2019), or deterministic neural networks to fit basis functions for Bayesian linear regression (Ma and Hernández-Lobato,, 2021) or the mean of a generalized sparse GP with Wasserstein-2 metric (Wild et al.,, 2022). Our work combines linearized BNNs with generalized variational inference, but we use the regularized KL divergence (Quang,, 2019), which naturally generalizes the conventional KL divergence and allows for informative GP priors.
Approximating function-space measures with weight-space priors in BNNs
Flam-Shepherd et al., (2017); Tran et al., (2022) minimize a divergence between the BNN’s prior predictive and a GP before performing inference on weights, while Wu et al., (2023) directly incorporate the bridging divergence inside the inference objective. Alternatively, Pearce et al., (2020) derive BNN architectures mirroring GPs, and Matsubara et al., (2022) use the Ridgelet transform to design weight-spaces priors approximating a GP in function space. Another line of work considers weight-space priors which regularize in function space by comparing the model’s predictions to those of a reference model (Nalisnick et al.,, 2020) and using an empirical prior (Rudner et al.,, 2023).
@@ 6 Discussion
We proposed a simple inference method with a well-defined variational objective function for Bayesian neural networks with Gaussian process (GP) priors in function-space. As standard VI with functions-space priors suffers from an infinite KL divergence problem, we propose to follow the generalized VI framework. Specifically, we substitute the conventional KL divergence in the ELBO by the regularized KL divergence, which is always finite, and which can be estimated consistently within the linearized BNN approximation. We demonstrated that our method allows to incorporate interpretable structural properties via a GP prior, accurately approximates the true GP posterior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
Future work should investigate the use of more expressive variational distributions, such as Gaussian with low-rank plus diagonal covariance proposed by Tomczak et al., (2020), which is compatible with our proposed method.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. Additional support was provided by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. Robert Bamler acknowledges funding by the German Research Foundation (DFG) for project 448588364 of the Emmy Noether Programme. The authors extend their gratitude to the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tristan Cinquin. Finally, we thank Marvin Pförtner for very useful discussions on the theory and experiments, and Vincent Fortuin for his feedback.
Appendix A Appendix
A.1 Divergences between Gaussian measures
A.1.1 The KL divergence is infinite
In this section, we show that the Kullbach-Liebler (KL) divergence between the Gaussian measures $\mathbb{Q}_{\phi}^{F}\sim\mathcal{N}\mathopen{}\left(m_{Q},C_{Q}\right)\mathclose{}$ and $\mathbb{P}^{F}\sim\mathcal{N}\mathopen{}\left(m_{P},C_{P}\right)\mathclose{}$, respectively induced by the linearized BNN in Eq 2.7 and by a non-degenerate Gaussian process satisfying conditions given in section 2.2, is infinite. While this has already been shown by Burt et al., (2020), the proof is easier for Gaussian measures. We first need the Feldman-Hàjek theorem which tells us when the KL divergence between two Gaussian measures is well-defined.
Theorem A.1 (Feldman-Hàjek, Quang, (2022) Theorem 2, Simpson, (2022) Theorem 7).
@@ A.1.1 The KL divergence is infinite
otherwise $\nu_{1}$ and $\nu_{2}$ are singular. If $\nu_{1}$ and $\nu_{2}$ are equivalent, then the Radon-Nikodym dervative exists and $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ admits an explicit formula. Otherwise, $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}=\infty$.
@@
A.1.2 The regularized KL divergence
In this section, we provide the bound describing the asymptotic convergence of the regularized KL divergence estimator.
Theorem A.2 (Convergence of estimator, Quang, (2022) Theorem 45).
@@ A.1.2 The regularized KL divergence
$|D_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(\mu_{1}(\mathbf{x}),K_{1}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}\,\|\,{\mathcal{N}\mathopen{}\left(\mu_{2}(\mathbf{x}),K_{2}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}}\right)\mathclose{}\\ -D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}\,\|\,\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}\right)\mathclose{}|\\ \leq\frac{1}{2\gamma}(B_{1}+B_{2})^{2}[1+\kappa_{2}^{2}/\gamma]\mathopen{}\left(\frac{2\log\frac{48}{\gamma}}{M}+\sqrt{\frac{2\log\frac{48}{\gamma}}{M}}\right)\mathclose{}\\ +\frac{1}{2\gamma^{2}}[\kappa_{1}^{4}+\kappa_{2}^{4}+\kappa_{1}^{2}\kappa_{2}^{2}(2+\kappa_{2}^{2}/\gamma)]\mathopen{}\left(\frac{2\log\frac{12}{\gamma}}{M}+\sqrt{\frac{2\log\frac{12}{\gamma}}{M}}\right)\mathclose{}$ (A.1)
A.2 Additional details on the GFSVI objective estimator
@@
$\mathcal{L}(\phi)=\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}$ (A.2)
@@
${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\mathcal{N}\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}}),\sigma_{y}^{2}\right)\mathclose{}\right]\mathclose{}=-\frac{1}{2}\log\mathopen{}\left(2\pi\sigma_{y}^{2}\right)\mathclose{}-\frac{\mathopen{}\left(y_{i}-f\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}\right)\mathclose{}^{2}+J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}{\bm{S}}J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}^{\top}}{2\sigma_{y}^{2}}.$ (A.3)
@@
$\begin{split}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\text{Cat}(y_{i}\,|\,\sigma(f_{L}({\bm{x}}_{i};{\bm{w}})))\right]\mathclose{}&=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathbb{I}[y_{i}=c]\Bigl{[}f^{c}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\\ &\qquad-\log\left[\sum_{c^{\prime}=1}^{C}\exp\left(f^{c^{\prime}}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\right)\right]\Bigr{]}\end{split}$ (A.4)
@@
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}=\frac{1}{2}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}^{\top}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}\\ +\frac{1}{2}\operatorname{Tr}\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})-\mathbb{I}_{M}\big{]}\\ -\frac{1}{2}\log\det\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})\big{]}$ (A.5)
@@
A.3 Additional details on the experimental setup
A.3.1 Experiments on synthetic data

### v3

@@ 4.1 Quantitative results on real-world data
Metric Function-space priors Weight-space priors GP-based
@@
We evaluate GFSVI on regression, classification, and out-of-distribution detection. In all tables, we bold the highest score and any score whose error bar (standard error) overlaps with the highest score’s error bar.
Table 3: Results for the ocean current modeling task.
Metric GFSVI (ours) TFSVI VIP GP
Log-like. -6.627 $\pm$ 0.753 -22.651 $\pm$ 2.947 -11.631 $\pm$ 3.171 -0.507 $\pm$ 0.000
MSE -0.021 $\pm$ 0.002 -0.034 $\pm$ 0.003 -0.026 $\pm$ 0.001 -0.013 $\pm$ 0.000
Ocean current modeling.
We measure how well GFSVI can incorporate knowledge specified via a GP prior on real-world data by considering the problem of modeling ocean currents in the Gulf of Mexico. We follow the setup by Shalashilin [2024] and use the GulfDrifters dataset [Lilly and Pérez-Brunius, 2021] to estimate ocean currents from $20$ $2$-dimensional velocity vectors collected from drifter buoys. We embed physical properties of fluid motions into the GP prior and to the neural networks by applying the Helmholtz decomposition [Berlinghieri et al., 2023, Cinquin et al., 2024]. We compare our GFSVI to a GP, to TFSVI and to VIP. More details can be found in section C.2. We find that incorporating knowledge via an informative GP prior in GFSVI improves performance over weight-space priors in TFSVI and VIP (see table 3 and figure 3). However, the GP outperforms both BNNs, which suggests that the physically motivated kernel describes the fluid dynamics well enough that the additional inductive bias introduced by a neural network hurts performance rather than helping it. In the following, we consider experiments with larger datasets (making exact GP inference computationally infeasible in many cases), and where structural prior knowledge in function space exists but is not derived from laws of nature.
@@
Regression.
We assess the predictive performance of GFSVI on data sets from the UCI repository [Dua and Graff, 2017]. Table 1, and Table 6 in the appendix, show expected log-likelihood and mean squared error, respectively. We perform 5-fold cross validation and report means and standard errors across the test folds. We also rank the methods for each dataset and report the mean rank of each method across all datasets. See section C.3 for more details. We find that GFSVI performs competitively compared to baselines and obtains the best mean rank for both metrics, matching the top performing methods on nearly all datasets. In particular, we find that using GP priors in the linearized BNN with GFSVI yields improvements over the weight-space priors used in TFSVI, and that GFSVI performs slightly better than FVI despite being simpler. Further, we find that GFSVI approximates the exact GP-posterior more accurately that FVI (see table 7 and section D.3), and that it converges in slightly more steps than TFSVI (figure 15).
Classification.
We further evaluate classification performance of our method on the MNIST [LeCun et al., 2010] and FashionMNIST [Xiao et al., 2017] image data sets. We fit the models on a random subset of 90% of the training set, use the remaining 10% as validation data, and evaluate on the provided test split. We repeat with 5 different random seeds and report the mean and standard error of the expected log-likelihood, accuracy, and expected calibration error (ECE) in table 2. For GFSVI, FVI, and TFSVI, we tested measurement points from both a uniform random (rnd) distribution $\rho({\bm{x}})$ and from kmnist. Details in section C.4. We find that GFSVI performs competitively on MNIST, exceeding the expected log-likelihood and accuracy of top-scoring baselines and similarly to best baselines on FashionMNIST. GFSVI also yields well-calibrated models with low ECE.
Out-of-distribution detection.
We next evaluate our method by testing if its epistemic uncertainty is predictive of out-of-distribution (OOD) data. We consider two settings: (i) with tabular data and a Gaussian likelihood [Malinin et al., 2021], and (ii) with image data and a categorical likelihood [Osawa et al., 2019]. We report the accuracy of classifying OOD vs. in-distribution (ID) data using a (learned) threshold on the predictive uncertainty. More details in section C.5. In setting (i), GFSVI performs competitively and obtains the highest mean rank (table 4). Likewise in setting (ii), GFSVI strongly outperforms all baselines when using the kmnist measurement point distribution $\rho({\bm{x}})$ (figure 12, tables 2 and 8). We find that with high-dimensional image data, the choice of measurement point distribution highly influences OOD detection accuracy (see Appendix D.5 for a discussion). In both settings, using GP priors with GFSVI rather than weight-space priors with TFSVI is beneficial, and GFSVI also improves over FVI. GFSVI’s uncertainty is also well-calibrated under distribution shift of the input features (see section D.6).
@@ 5 Related work
In this section, we review related work on function-space VI with neural networks, and on approximating functions-space measures with weight-space priors.
Function-space inference with neural networks.
Prior work on function-space VI in BNNs has addressed issues (i) intractable variational posterior in function space and (ii) intractable KL divergence discussed in Section 2.1. Sun et al. [2019] address (i) by using implicit score function estimators, and (ii) by replacing the supremum with an expectation. Rudner et al. [2022b] address (i) by using a linearized BNN [Khan et al., 2019, Immer et al., 2021, Maddox et al., 2021], and (ii) by replacing the supremum with a maximum over a finite set. Other work abandons approximating the neural network’s posterior and instead uses a BNN to specify a prior [Ma et al., 2019], or deterministic neural networks as features for Bayesian linear regression [Ma and Hernández-Lobato, 2021] or the mean of a generalized sparse GP [Wild et al., 2022b]. Unlike our more expressive GP posterior covariance, Wild et al. [2022b] uses a simple stationary sparse GP posterior covariance (table 2) which has higher sampling cost and can lead to model misspecification (figure 14). Our work combines linearized BNNs with generalized VI, but we use the regularized KL divergence [Quang, 2019], which naturally generalizes the KL divergence and allows for informative GP priors.
Approximating function-space measures with weight-space priors.
Flam-Shepherd et al. [2017], Tran et al. [2022] minimize a divergence between the BNN’s prior predictive and a GP before performing inference on weights, while Wu et al. [2023] directly incorporate the bridging divergence inside the inference objective. Alternatively, Pearce et al. [2020] derive BNN architectures mirroring GPs, and Matsubara et al. [2021] use the Ridgelet transform to design weight-spaces priors approximating a GP in function space. Similarly, Rudner et al. [2023] and Sam et al. [2024] use empirical weight-space priors to regularize in function space and encode domain knowledge specified via a loss function, respectively. Yang et al. [2020] instead imposes functional constraints directly via the prior.
@@ 6 Discussion
We proposed a simple inference method with a well-defined variational objective function for BNNs with GP priors in function-space. As standard VI with functions-space priors suffers from an infinite KL divergence problem, we propose to follow the generalized VI framework. Specifically, we substitute the conventional KL divergence in the ELBO by the regularized KL divergence, which is always finite, and which can be estimated consistently within the linearized BNN approximation. We demonstrated that our method allows to incorporate interpretable structural properties via a GP prior, accurately approximates the true GP posterior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNNs with both function-space and weight-space priors. Future work should investigate the use of more expressive variational distributions, such as Gaussian with low-rank plus diagonal covariance proposed by Tomczak et al. [2020].
Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645, and by the German Research Foundation (DFG) under project 448588364 of the Emmy Noether Programme. Additional support was provided by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. The authors extend their gratitude to the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tristan Cinquin. Finally, Tristan Cinquin thanks Marvin Pförtner for very useful discussions on the theory and experiments, and Vincent Fortuin for feedback.
Supplementary Material
Appendix A Divergences between Gaussian measures
A.1 The KL divergence is infinite
In this section, we show that the Kullbach-Liebler (KL) divergence between the Gaussian measures $\mathbb{Q}_{\phi}^{F}\sim\mathcal{N}\mathopen{}\left(m_{Q},C_{Q}\right)\mathclose{}$ and $\mathbb{P}^{F}\sim\mathcal{N}\mathopen{}\left(m_{P},C_{P}\right)\mathclose{}$, respectively induced by the linearized BNN in Eq 2.7 and by a non-degenerate Gaussian process satisfying conditions given in section 2.2, is infinite. While this has already been shown by Wild et al. [2022a], the proof is easier for Gaussian measures. We first need the Feldman-Hàjek theorem which tells us when the KL divergence between two Gaussian measures is well-defined.
Theorem A.1 (Feldman-Hàjek, Quang [2022] Theorem 2, Simpson [2022] Theorem 7).
@@ A.1 The KL divergence is infinite
otherwise $\nu_{1}$ and $\nu_{2}$ are singular. If $\nu_{1}$ and $\nu_{2}$ are equivalent, then the Radon-Nikodym derivative exists and $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ admits an explicit formula. Otherwise, $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}=\infty$.
@@
A.2 The regularized KL divergence
We provide the bound describing the asymptotic convergence of the regularized KL divergence estimator in equation A.1. The error results from the fact that taking a finite number M of context points effectively cuts off the spectra of the covariance operators and the estimator $\hat{D}_{\text{KL}}^{\gamma}$ converges to $D_{\text{KL}}^{\gamma}$ as $M\to\infty$ with high probability.
Theorem A.2 (Convergence of estimator, Quang [2022] Theorem 45).
@@ A.2 The regularized KL divergence
$|D_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(\mu_{1}(\mathbf{x}),K_{1}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}\,\|\,{\mathcal{N}\mathopen{}\left(\mu_{2}(\mathbf{x}),K_{2}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}}\right)\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}\,\|\,\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}\right)\mathclose{}|\\ \leq\frac{1}{2\gamma}(B_{1}+B_{2})^{2}[1+\kappa_{2}^{2}/\gamma]^{2}\mathopen{}\left(\frac{2\log\frac{48}{\delta}}{M}+\sqrt{\frac{2\log\frac{48}{\delta}}{M}}\right)\mathclose{}\\ +\frac{1}{2\gamma^{2}}[\kappa_{1}^{4}+\kappa_{2}^{4}+\kappa_{1}^{2}\kappa_{2}^{2}(2+\kappa_{2}^{2}/\gamma)]\mathopen{}\left(\frac{2\log\frac{12}{\delta}}{M}+\sqrt{\frac{2\log\frac{12}{\delta}}{M}}\right)\mathclose{}$ (A.1)
Note that equation A.1 provides a very general bound on the error that does not make assumptions on the spectral decay, and it may therefore dramatically overestimate the error. Indeed, we analyze convergence empirically in figure 20 and observe that the estimator converges quickly except for very rough priors (e.g., Matérn-1/2) with very small $\gamma$.
Appendix B Additional details on the GFSVI objective estimator
@@
$\mathcal{L}(\phi)=\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}$ (B.1)
@@
${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\mathcal{N}\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}}),\sigma_{y}^{2}\right)\mathclose{}\right]\mathclose{}=-\frac{1}{2}\log\mathopen{}\left(2\pi\sigma_{y}^{2}\right)\mathclose{}-\frac{\mathopen{}\left(y_{i}-f\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}\right)\mathclose{}^{2}+J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}{\bm{S}}J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}^{\top}}{2\sigma_{y}^{2}}.$ (B.2)
@@
${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\text{Cat}(y_{i}\,|\,\sigma(f_{L}({\bm{x}}_{i};{\bm{w}})))\right]\mathclose{}=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathbb{I}[y_{i}=c]\left[f^{c}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})-\log\left[\sum_{c^{\prime}=1}^{C}\exp\left(f^{c^{\prime}}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\right)\right]\right]$ (B.3)
@@
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}=\frac{1}{2}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}^{\top}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}\\ +\frac{1}{2}\operatorname{Tr}\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})-\mathbb{I}_{M}\big{]}\\ -\frac{1}{2}\log\det\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})\big{]}$ (B.4)
@@
Computational complexity
Evaluating the objective in Eq. B.1 has complexity $O(BKC+M^{3})$ for Categorical likelihoods and $O(B+M^{3})$ for Gaussian likelihoods, where $B$ is the batch size, $K$ the number of variational posterior samples, $C$ the number of classes, and $M$ the number of context points. The first term corresponds to the expected log-likelihood in our objective and the second term to the regularized KL divergence estimator. We note that evaluating the linearized neural network can be efficiently done in about 3x the cost of one forward pass using the Jacobian-vector product computational primitive.
Appendix C Additional details on the experimental setup
C.1 Experiments on synthetic data